The results of the nationwide aptitude test in mathematics are normally distributed with a mean of 80 and a standard deviation of 15. Find the raw score such that 70% of the cases are below it.




1
Expert's answer
2022-04-03T15:25:07-0400

"P(Z<z)=0.7 \to z=0.52."

"\frac{x-80}{15}=0.52 \to x=0.52*15+80=88."